OpenStudy (studygurl14):

Hello! Welcome to StudyGurl14's Study Session! Feel free to ask any questions and I'll try my best to walk you through it and give you an answer. :) Satisfaction guaranteed! (Maybe...lol) Medals are appreciated if you receive the answer you are looking for, but not in any way required. Thanks and let the questions begin! ~StudyGurl14

OpenStudy (anonymous):

For each figure, find the value of x. x = _____° heres the picture: http://takeexam.next.ecollege.com/ec/pub/K12Content/Courses/resources/swfs/media/Ninth_Grade/Math/Unit6664/Lesson38095/VHS_GMT_03_10_quiz1.gif

OpenStudy (studygurl14):

I got an error message when i clicked the link...

OpenStudy (mathmath333):

\(\huge\tt \color{black}{\sum\limits_{n=1}^{100}n.2^n}\) i need formula for this series

OpenStudy (studygurl14):

I don't understand what you need @mathmath333, could you explain? What's the question say?

OpenStudy (mathmath333):

\(\large\tt \color{black}{1.2+2.2^2+3.2^3....100.2^{100}}\)

OpenStudy (studygurl14):

It looks like you already have the formula there...so I'm confused still

OpenStudy (studygurl14):

Sorry

OpenStudy (mathmath333):

i wanted to know how u derive the formula

OpenStudy (mathmath333):

its ok

OpenStudy (studygurl14):

Oh! You want to know how they got that formula?!

OpenStudy (studygurl14):

Is that what you want?

OpenStudy (mathmath333):

yes!

OpenStudy (studygurl14):

Okay. I get it now lol

OpenStudy (studygurl14):

Okay, so that sign is called sigma notation. Does that sound familiar?

OpenStudy (mathmath333):

yes

OpenStudy (studygurl14):

are the periods supposed to be multiplication signs...?

OpenStudy (studygurl14):

or are they decimals?

OpenStudy (studygurl14):

is 1.2 supposed to mean 1.2 or 1•2?

OpenStudy (mathmath333):

\(\huge\tt \color{black}{\times}\)

OpenStudy (mathmath333):

i hope u got my question

OpenStudy (studygurl14):

Okay so the n=1 means you substitute 1 for n in the equation \[n \times2^n\]

OpenStudy (mathmath333):

i have the formula for nth term and i need the summation formula upto nth term for it

OpenStudy (studygurl14):

and then, to get the next number in the sequence, you plug in what you got for \[(1) \times 2^1\] as n into the equation \[n \times 2^n\]...and so forth

OpenStudy (studygurl14):

the 100 indicates the amount of terms you are including in the summation. Does what I've said so far make sense?

OpenStudy (mathmath333):

u know formula for first n natural numbers like that i need formula for this series i hope u understand my question

OpenStudy (studygurl14):

...like finding the nth term of a sequence?

OpenStudy (mathmath333):

no sum upto nth term of sequence

OpenStudy (studygurl14):

http://www.purplemath.com/modules/series4.htm look at example #2

OpenStudy (anonymous):

In triangle RST, measure of angle R = (2x + 10), measure of angle S = (2x + 25), and measure of angle T = (x - 5). Find the value of x.

OpenStudy (studygurl14):

All triangles add up to 180 degrees, so... (2x + 10) + (2x + 25) + (x - 5) = 180

OpenStudy (studygurl14):

Combine like terms: 5x + 30 = 180 5x = 150 x = 30

OpenStudy (anonymous):

Great thank you for explaining it!!

OpenStudy (haichi):

Acid or base

OpenStudy (anonymous):

Suppose the measure of an exterior of an equiangular triangle is (x + 15). What is the value of x?

OpenStudy (studygurl14):

yw @madmanty

OpenStudy (anonymous):

yw?

OpenStudy (studygurl14):

yw = you're welcome

OpenStudy (anonymous):

oh haha ok

OpenStudy (studygurl14):

interior angle + exterior angle = 180 degrees equangular triangle means all angles equal 60 degrees so 60 + ? = 180

OpenStudy (studygurl14):

60 + (x + 15) = 180

OpenStudy (studygurl14):

x=?

OpenStudy (anonymous):

ohhhhh ok. thanks again!! haha x=105?

OpenStudy (studygurl14):

correct!

OpenStudy (anonymous):

sweet!! thanks so much!

OpenStudy (anonymous):

the compliment of an angle measures 42 degrees. what is the measure of the angle?

OpenStudy (studygurl14):

complimetnary angles add up to 90 degrees' 90 - 42 = answer

OpenStudy (anonymous):

your a genius aha

OpenStudy (anonymous):

What are converse statements?

OpenStudy (anonymous):

what is the measure of angle x?

OpenStudy (studygurl14):

Converse statemtns: http://hotmath.com/hotmath_help/topics/converse-inverse-contrapositive.html

OpenStudy (anonymous):

i figured that one out, but i need help with the last one i asked..

OpenStudy (studygurl14):

use your knowledge of exterior/interior angles and supplementary angles